# Opponent Shaping for Antibody Development

### v2

@@ 2.3.1 Attack is the Best Defence
Figure 3: Effect of myopic and shaper antibodies on viral antigens induced by other antibodies. We optimise $80$ different antibodies for each horizon (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps and we group the escaped viruses $v_{H}$ by the horizon $H$ of the antibody that induced them, these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies against the final escape viral variant $v_{H}$ induced by other antibodies $a_{H^{\prime}}$. Darker colours correspond to better antibody payoff.
To discern which strategy our antibody shapers adopt, we investigate whether shapers optimised with different horizons exert different types of evolutionary pressure on the virus. We generate antibodies for each horizon (Myopic, $H=5,H=10,H=20,H=100$) and simulate viral escape against each antibody for 100 steps, resulting in viruses $v_{H}$. We then evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each horizon antibody against the final viral variants $v_{H}$ induced by other antibodies $a_{H^{\prime}}$. Figure 3 presents the result of this analysis.
Interestingly, viruses $v_{100}$ induced by $H=100$ shapers are consistently more exploitable by antibodies across all optimisation horizons. This suggests that $H=100$ shapers actively shape the escape trajectories of the virus in a way that makes the resulting variants more susceptible to antibody binding in general.
However, this shaping effect comes at a cost. The $a_{100}$ shapers show slightly lower payoffs compared to the peak performance of shorter-horizon antibodies ($a_{5}$ and $a_{10}$) against the viruses $v_{100}$ induced by the $\mathit{H=100}$ shapers (see right most column of Figure 3). This trade-off indicates that to exert a stronger shaping influence on viral evolution, $\mathit{H=100}$ shapers sacrifice some degree of immediate performance or robustness, that is their ability to perform well against a wide range of viruses. Therefore, a potential strategy could involve using a “mixture” of antibodies as therapy: some optimised for shaping the virus’s evolutionary trajectory, and others designed for strong immediate binding.
@@ 2.3.2 Amino Acid Distribution in Shapers and Myopic Antibodies
To further understand the performance differences between myopic antibodies and shapers, we analyse the amino acid distributions of the antibodies. We compared myopic antibodies with shapers optimised using horizons $H={5,10,20,100}$ over $N=30$ antibody optimisation steps. Figure 4 showcases the results of our experiment.
@@
In contrast, the clustering behaviour we observe in shorter-horizon antibodies indicates a more specialised strategy. By concentrating on amino acids at the extremes of the binding energy spectrum, these antibodies may achieve strong immediate binding but potentially at the cost of long-term robustness.
Figure 4: Distribution of amino acids in myopic antibodies and shapers. The antibodies are optimised using different viral escape horizons $H$. Longer horizon shapers push the amino acid distribution closer to a uniform distribution.
However, while this analysis hints at the robustness of long-term shapers, it does not fully explain the shaping behaviour we observed in our previous results. In the next section, we investigate the distribution of amino acids within specific binding poses.
@@ 2.3.3 Influence of Antibody Shapers on Binding Poses
As both the virus and the antibody mutate during our optimisation process, the lowest energy pose can change. To capture these dynamics, we introduced the concept of a pose matrix: a $20\times 20$ matrix11120202020 is the number of possible amino acids. where one dimension corresponds to the antibody amino acids and the other to the viral amino acids, both contributing to the lowest energy pose. The entries in this matrix represent the number of interactions between specific amino acid pairs.
Figure 5a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe two key trends:
  1. 1.
    As viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape.
  2. 2.
    As the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure 5b) and the virus axis (Figure 5c). These figures show the change in interaction counts between viral escape steps 00 and $100$.
Figure 5b reveals a viral escape strategy: including more of the antibody’s lowest binding amino acids (particularly K) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K interactions. Furthermore, Figure 5c shows another viral escape strategy: removing its high-binding amino acids (I and M) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.
Based on these observations, we hypothesise that the shaping ability of $H=100$ shapers relies on two main mechanisms:
  1. 1.
    Preventing the virus from including the antibody’s lowest binding amino acids in the pose.
  2. 2.
    Inhibiting the virus from removing its own high-binding amino acids from the pose.
These strategies constrain the viral escape trajectories, making the resulting viral variants more susceptible to antibody binding in general.
While these results are specific to our Absolut! binding simulator, they demonstrate that the behaviour of antibody shapers is both explainable and intuitive. This work serves as a proof of concept, showing that opponent shaping techniques can optimise long-horizon shapers to produce antibodies that more effectively prevent viral escape.
The transferability of these findings to more realistic binding simulators and viral escape models remains an open question. However, these results are encouraging: we have demonstrated that shapers not only perform better and shape viral escape trajectories, but we can also explain the strategies they employ to achieve this shaping effect. This analysis underscores the potential of opponent shaping in antibody design and opens new avenues for developing more effective, long-term strategies against rapidly evolving pathogens.
@@ 3 Discussion
Furthermore, our analysis reveals that long-horizon shapers shape the virus into more exploitable variants as opposed to just becoming robust to any virus, see Figure 3. In fact, the short-horizon shapers bind stronger to the viruses included by the long-horizon shapers than the long-horizon shapers themselves. This suggests exploring a mixture of antibodies, some optimised for virus shaping and others for binding. This could leverage the benefits of both shaping the viral evolutionary landscape and maintaining strong binding efficacy. Interestingly, a longitudinal study on rare patients who naturally developed broadly neutralizing antibodies (bnAbs) to HIV showed that early neutralizing antibodies induce HIV escape to those antibodies [22]. The bnAbs were later developed with the cooperation between the B cell lineages. This indicates the constant evolving dynamics between the virus and our immune system in nature. In the future, our approach could enable us to better understand and develop therapies for similar settings by simulating a game with more than one antibody playing against the virus.
To provide insights into the mechanisms underlying the effectiveness of antibody shapers, we conduct an explainability analysis focused on the amino acid distribution in shapers (Figure 4) and on binding poses (Figure 5). While these results offer insights into the strategies employed by our antibody shapers, it’s important to note that they are specific to the Absolut! binding model [12, 14]. Antibodies designed using more accurate binding simulators might be able to employ other strategies in their interactions with viruses. Nevertheless, the fact that we can explain the behaviour of our antibody shapers is encouraging. It suggests that, even as we move towards more complex and realistic models, we may be able to maintain a degree of interpretability in our approach.
While our approach shows promise, it’s important to acknowledge its current limitations, particularly in the modelling of antibody-antigen interactions. Our study relies on the Absolut! framework which is a dramatically simplified model of reality.[12] . Moreover, here we focus only on the CDRH3 binding domain of the antibody and consider a fixed length (11 amino acid) region. Insertions and deletions are not within our scope of the study. Regardless of the CDRH3 region being the most critical for both antibody specificity and affinity [15], the other five binding regions (CDR1-3 on both the heavy and light chains) also partially contribute to the binding profiles, either by making direct contacts with the antigen or putting constraints on the conformational space. In future work, we want to extend the binding simulation to all the CDR regions with both light and heavy chains.
@@
Another promising avenue for enhancing our framework lies in the integration of more accurate viral escape forecasting methods such as EVEscape [24]. EVEscape successfully anticipated the emergence of COVID-19 variants using only pre-pandemic data. Incorporating such sophisticated viral escape forecasting as the simulated viral escape component of our framework could significantly boost its real-world applicability. This integration can help to bridge the gap between our computational approach and practical therapeutic development.
@@ 4.4 Binding Function
To give a more concrete example, for this paper, we use the dengue virus antigen [21]. Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation 1000100010001000444In practice, the difference is closer to $10,000$ times faster for a negligible performance drop for this particular antigen sequence.
@@
Using these methods of pruning poses gives us two subsets of the original set of poses, a larger one which almost exactly matches performance, and another which sometimes differs, but is much faster to compute. We refer to these as the high-resolution and low-resolution binding simulators respectively.
Note that for the low-resolution binding simulator, the more mutations the virus undergoes the less accurate it becomes. Furthermore, we also do binding to the antibody anti-target, $t^{-}_{a}$. To account for this, we compute the relevant poses for this anti-target too.

### v4

@@ 2.3.1 Attack is the Best Defence
Figure 3: Effect of myopic and shaper antibodies on viruses induced by other antibodies. We optimise $80$ different antibodies $a_{H}$ for each horizon (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps and we group the escaped viruses $v_{H}$ by the horizon $H$ of the antibody that induced them, these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies $a_{H^{\prime}}$ against the final escape viral variant $v_{H}$ induced by other antibodies optimised with horizon $H$. Darker colours correspond to better antibody payoff.
To disentangle these strategies, we seperately evaluate the antibodies and the viruses that evolve in response to them. To do this, we compare the virus against other antibodies which did not influence the viral evolution. The intuition is that an antibody which is good at shaping (good attack), but less robust (poor defence), will induce viruses which other antibodies will perform well against. Specifically, we generate antibodies $a_{H}$ for each horizon $H$ and simulate viral escape against these antibodies for 100 steps, resulting in viruses $v_{H}$. For all pairs of horizons $(H,H^{\prime})$, we then cross-evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$. Figure 3 presents the result of this analysis.
Interestingly, viruses $v_{100}$ induced by $H=100$ shapers are consistently more exploitable by antibodies across all optimisation horizons. This suggests that $H=100$ shapers actively shape the escape trajectories of the virus in a way that makes the resulting variants more susceptible to antibody binding in general. However, this shaping effect comes at a cost. The $H=100$ shapers ($a_{100}$) show slightly lower payoffs compared to the peak performance of shorter-horizon antibodies ($a_{5}$ and $a_{10}$) against the viruses $v_{100}$ induced by the $H=100$ shapers (see rightmost column of Figure 3). This trade-off indicates that to exert a stronger shaping influence on viral evolution, $H=100$ shapers sacrifice some degree of immediate performance or robustness, that is their ability to perform well against a wide range of viruses. Therefore, a potential strategy could involve using a “mixture” of antibodies as therapy: some optimised for shaping the virus’s evolutionary trajectory, and others designed for strong immediate binding.
@@ 2.3.2 Amino Acid Distribution in Shapers and Myopic Antibodies
To further understand the performance differences between myopic antibodies and shapers, we analyse how the amino acid distributions of the antibodies change with the optimisation horizon $H$. Within our model, each amino acid is solely characterised by its binding strength to other amino acids as defined by the Miyazawa-Jernigan energy potential matrix [14]. However, despite this simplicity, we still see interesting patterns in the amino acid distribution. Figure 4 showcases the results of our experiment.
@@
In contrast, the clustering behaviour we observe in shorter-horizon antibodies indicates a more specialised strategy. By concentrating on amino acids at the extremes of the binding energy spectrum, these antibodies may achieve strong immediate binding but potentially at the cost of long-term robustness. However, while this analysis hints at the robustness of long-term shapers, it does not fully explain the shaping behaviour we observed in our previous results. In the next section, we investigate the distribution of amino acids within specific binding poses.
Figure 4: Distribution of amino acids in myopic antibodies and shapers. The antibodies are optimised for $N=30$ steps using different viral escape horizons $H$. Longer horizon shapers push the amino acid distribution closer to a uniform distribution.
@@ 2.3.3 Influence of Antibody Shapers on Binding Poses
As both the virus and the antibody mutate during our optimisation process, the lowest energy pose can change. To capture these dynamics, we introduced the concept of a pose matrix: a $20\times 20$ matrix22220202020 is the number of possible amino acids. where one dimension corresponds to the antibody amino acids and the other to the viral amino acids, both contributing to the lowest energy pose. The entries in this matrix represent the number of interactions between specific amino acid pairs.
Figure 5a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe the following two trends. We observe two key trends. First, as viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape. Second, as the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure 5b) and the virus axis (Figure 5c). These figures show the change in interaction counts between viral escape steps 00 and $100$. Figure 5b shows that as the virus escapes it includes more of the antibody’s lowest binding amino acids (particularly K) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K interactions. Furthermore, Figure 5c shows another viral escape strategy, where the virus removes its high-binding amino acids (I and M) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.
Based on these observations, we hypothesise that the shaping ability of $H=100$ shapers relies on two main mechanisms. Preventing the virus from including the antibody’s lowest binding amino acids in the pose, and inhibiting the virus from removing its own high-binding amino acids from the pose. These strategies constrain the viral escape trajectories, making the resulting viral variants more susceptible to antibody binding in general. While these results are specific to our Absolut! binding simulator, they demonstrate that the behaviour of antibody shapers is both explainable and intuitive. This work serves as a proof of concept, showing that opponent shaping techniques can optimise antibodies to more effectively prevent viral escape.
@@ 3 Discussion
Furthermore, our analysis reveals that long-horizon shapers shape the virus into more exploitable variants as opposed to just becoming robust to any virus, see Figure 3. In fact, the short-horizon shapers bind stronger to the viruses included by the long-horizon shapers than the long-horizon shapers themselves. This suggests exploring a mixture of antibodies, some optimised for virus shaping and others for binding. This could leverage the benefits of both shaping the viral evolutionary landscape and maintaining strong binding efficacy. Interestingly we observe a similar “strategy” deployed by the human body: a longitudinal study on rare patients who naturally developed broadly neutralizing antibodies (bnAbs) to HIV showed that early neutralizing antibodies induce HIV escape to those antibodies [23]. In the future, our approach could enable us to better understand and develop therapies for similar settings by simulating a game with more than one antibody playing against the virus.
To provide insights into the mechanisms underlying the effectiveness of antibody shapers, we conduct an explainability analysis focused on the amino acid distribution in shapers (Figure 4) and on binding poses (Figure 5). While these results offer insights into the strategies employed by our antibody shapers, it’s important to note that they are specific to the Absolut! binding model [11, 14], and the transferability of our findings to more realistic binding simulators and viral escape models remains an open question. Antibodies designed using more accurate binding simulators might be able to employ other strategies in their interactions with viruses, and the strategies we explore in this model may not apply in a more realistic simulation. Nevertheless, the fact that we can explain the behaviour of our antibody shapers is encouraging. It suggests that, even as we move towards more complex and realistic models, we may be able to maintain a degree of interpretability in our approach.
@@
A consequence of us using the Absolut! framework [11] is that we make certain modelling assumptions. Namely, we focus only on the CDRH3 binding domain of the antibody and consider a fixed length (11 amino acid) region. Regardless of the CDRH3 region being the most critical for both antibody specificity and affinity [15], the other five binding regions (CDR1-3 on both the heavy and light chains) also partially contribute to the binding profiles, either by making direct contacts with the antigen or putting constraints on the conformational space. In future work, we want to extend the binding simulation to all the CDR regions with both light and heavy chains.
Another promising avenue for enhancing our framework lies in the integration of more accurate viral escape forecasting methods [25, 26]. EVEscape [25] successfully anticipated the emergence of COVID-19 variants using only pre-pandemic data. Incorporating such sophisticated viral escape forecasting as the simulated viral escape component of our framework could significantly boost its real-world applicability. This integration can help to bridge the gap between our computational approach and practical therapeutic development.
@@ 4.4 Binding Function
To give a more concrete example, for this paper, we use the dengue virus antigen [22]. Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation 1000100010001000555In practice, the difference is closer to $10,000$ times faster for a negligible performance drop for this particular antigen sequence.
@@
Using these methods of pruning poses gives us two subsets of the original set of poses, a larger one which almost exactly matches performance, and another which sometimes differs, but is much faster to compute. We refer to these as the high-resolution and low-resolution binding simulators respectively. Note that for the low-resolution binding simulator, the more mutations the virus undergoes the less accurate it becomes. Furthermore, we also do binding to the antibody anti-target, $t^{-}_{a}$. To account for this, we compute the relevant poses for this anti-target too.
@@
5 Data Availability
For our experiment we use the envelope protein E from the dengue viral antigen, PDB code 2R29 [22]. The structure data is available in the PDB [21] under https://www.rcsb.org/structure/2R29. The Absolut! [11] binding pose data we use for our JAX [13] implementation is all available under https://shorturl.at/NuUfN.
6 Code Availability
All code we use to implement our method and obtain the results presented here is available in the GitHub repository https://github.com/olakalisz/antibody-shapers.